OpenStudy (anonymous):

When solving this equation, y = x² + 6x + 5, compare the method of factoring to the quadratic formula to find the roots. Which way seems easier and do you think it will always be this way?

OpenStudy (anonymous):

lets see factoring methid.. can you think of two numbers a, b such trhat a times b = 5 a plus b = 6

OpenStudy (anonymous):

5x5=1? and 3+3=6?

OpenStudy (anonymous):

5 times 5 = 25

OpenStudy (anonymous):

sorry I meant 5x1

OpenStudy (anonymous):

lets say the numbers are 1 and 5 then 1 times 5 = 5 and 1 plus 5 = 6

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

and this satisfies our quadratic. remember, you have to tick with two same numbers \[y = x^2+1x+5x+5\\ y = x(x+1)+5(x+1)\\ y=(x+1)(x+5) \] so, we were easily able to factorize it

OpenStudy (anonymous):

now, using the quadratic formula, \[x={-b\pm\sqrt{b^2-4ac}\over2a}\\ a=1\quad b=6\quad c=5\\ x=\frac{-6\pm\sqrt{36-4(1)(5)}}{2(1)}\\ x=\frac{-6\pm\sqrt{16}}{2}\\ x={-6\pm4\over2}\\ x={-6+4\over2}\qquad x={-6-4\over2}\\ x=-1\qquad x=-5 \]

OpenStudy (anonymous):

which one is easier?

OpenStudy (anonymous):

I would say factoring

OpenStudy (anonymous):

good. but can you always use that?

OpenStudy (anonymous):

no?

OpenStudy (anonymous):

good and why not?

OpenStudy (anonymous):

because you may not always be able to find two numbers that satisfy the consitions for example: \[y=x^2+x+1\] can you factorize this

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[y=x^2+2x+1\]can you factorize this?